Find the average rate of change for f(x)= 4x² from 0 to 7/4
and the equation of the secant line containing (0, f(0)) and (7/4, f(7/4)). Please show all of your work.

Answer by stanbon(75887)

Find the average rate of change for f(x)= 4x² from 0 to 7/4
and the equation of the secant line containing (0, f(0)) and (7/4, f(7/4)). Please show
----------------
The average rate of change is the slope of the line joining the two points.
---
slope = [f(7/4)-f(0)]/[(7/4)-0]
----
f(7/4) = 4(7/4)^2 = 49/4
------
slope = [(49/4)-0]/[7/4] = (49/4)/[7/4] = 7
=========================================
Equation of the line thru (0,0) and with slope = 7
---
Equation Form: y = mx + b
0 = 7*0 + b
b = 0
---
Equation of the secant: y = 7x
==================================
Cheers,
Stan H.
==========